OpenStudy (superteenblue):

A garden pond is 6 feet wide. The perimeter of the pond is 28 feet. What is the area of the pond?

OpenStudy (ellalaynemaguire):

48 feet squared

OpenStudy (anonymous):

what is shape of pond?

OpenStudy (tommo_lcfc):

We know one of the lengths of the pond is 6 feet. If the perimeter is 28 feet then the other length is 8 feet. Therefore the area of the pond is 6 x 8 = 48 ft^2

OpenStudy (superteenblue):

what?

OpenStudy (superteenblue):

tmmo hlp plz what do you mean

OpenStudy (tommo_lcfc):

If the pond is 6 ft wide and the perimeter of the pond is 28 square ft, then what is the length?

rebeccaxhawaii (rebeccaxhawaii):

i assume its a rectangle pond - so p=L+L+W+W so 28-6-6=1 then 16/2=your length

rebeccaxhawaii (rebeccaxhawaii):

so then to get area A=LW so A=8*6

OpenStudy (tommo_lcfc):

What she said!

OpenStudy (superteenblue):

im a fourth grader im not in high school i don understand

OpenStudy (superteenblue):

hlp plz

rebeccaxhawaii (rebeccaxhawaii):

okay so your perimeter is 28 ft

rebeccaxhawaii (rebeccaxhawaii):

the width is 6 ft the formula to get perimeter is length+length+width+width

OpenStudy (superteenblue):

ok so... whats the answer

rebeccaxhawaii (rebeccaxhawaii):

so length+length+6+6=28

rebeccaxhawaii (rebeccaxhawaii):

what is 6+6=?

OpenStudy (superteenblue):

so whats the area

OpenStudy (ellalaynemaguire):

if one side is 6 feet then you have to multiply it by 2 to represent the other side. then you subtract 12 from 28 because 6 times 2 is 12. 28-12 is 16 so you would divide that in half to find the length of the other 2 sides. 16 divided by 2 is 8. now that you have your width and length measurements you have to multiply them to find the area. 6 times 8 is 48 so your area would be 48 feet squared.

OpenStudy (ellalaynemaguire):

do you understand now?

OpenStudy (superteenblue):

i guess..